# No Free Lunch from Random Feature Ensembles

### v1

@@ 5 Width-Bottlenecked Scaling of Random Feature Ensembles
We now study the behavior of RFRR ensembles in the under-parameterized regime where $N=M/K\ll P$. Here, model size $N$ is the “bottleneck” that dictates the scaling behavior of the risk (Bahri et al., 2024; Maloney et al., 2022; Bordelon et al., 2024a; Atanasov et al., 2024). In particular, we ask how $E_{g}^{K}$ scales with total model size $M$ when $K$ and $N$ scale jointly with $M$ according to a “growth exponent” $\ell$:
@@
so that when $\ell=0$ the ensemble grows with $M$ by adding additional ensemble members of a fixed size, and when $\ell=1$ the ensemble grows by adding parameters to a fixed number of networks. Under the standard “source” and “capacity” constraints on the task eigenstructure (Cui et al., 2023; Caponnetto & De Vito, 2006; Bordelon et al., 2020; Defilippis et al., 2024) (the kernel eigenspectrum decays as $\eta_{t}\sim t^{-\alpha}$ with $\alpha>1$ and the target weights decay as $\bar{w}_{t}^{2}\eta_{t}\sim t^{-(1+2\alpha r)}$), we find:
$E_{g}^{K}\sim M^{-s},\qquad s=\min\left(2\alpha\ell\min(r,1),1-\ell+2\alpha\ell\min\left(r,\frac{1}{2}\right)\right)\,,$ (22)
with $s=2\alpha\ell\min(r,1)$ corresponding to the scaling of the bias and $s=1-\ell+2\alpha\ell\min\left(r,1/2\right)$ corresponding the scaling of the variance (reduced by a factor of $1/K$). A full derivation is provided in appendix C. These results reify the “no free lunch” result, as the optimal scaling law is always achieved when $\ell=1$. For difficult tasks, defined as having $r<1/2$, bias always dominates the error scaling and the scaling exponent increases linearly with $\ell$. However, when $r>1/2$, there will be a certain value $\ell^{*}$ above which error scaling is dominated by the variance term. When $r>1/2$, the scaling exponent of the variance increases from $1$ to $\alpha$ over the range $\ell\in[0,1]$. If $\alpha\gtrsim 1$, this can approach a flat line, and the dependence of the scaling exponent on $\ell$ can become weak, so that near-optimal scaling can be achieved for any $\ell>\ell^{*}$. When $11/2<r<$ 1, this transition occurs at $\ell^{*}=1/(1+\alpha(2r-1))$ and when $r>1$ it occurs at $\ell^{*}=1/(1+\alpha)$.
We plot these scaling laws with $\alpha=1.5$ in the regimes where $r<1/2$, $11/2<r<$ 1, and $r>1$ in fig. 3, along with the results of numerical simulations of linear RF regression on synthetic Gaussian datasets. As anticipated, for difficult tasks, where $r<1/2$, the scaling law improves linearly with $\ell$. However, for easier tasks ($r>1/2$), we see that, a near-optimal scaling law can be achieved as long as $\ell>\ell^{*}$ (fig. 3, center and right columns).
@@
Figure 4: Scaling laws provide a good description of width-bottlenecked RFRR ensembles.(A) we plot error as a function of $M$ at optimal ridge value for $\operatorname{ReLU}$ random-feature models applied to the binarized CIFAR-10 (left) and MNIST (right) classification tasks. (B) We plot theoretically predicted scaling exponents (eq. 22) for the bias and variance contributions to risk, as well as empirical power-law fits to risk in numerical simulations of RFRR models (see Appendix D.3, fig. S7)
6 No Free Lunch From Feature Learning Ensembles
Theorem 2 shows that for linear models, any benefit to ensembling under a fixed parameter count is a pathology of over-fitting. Provided a proper regularization is used to prevent over-fitting, a single large model will achieve optimal performance. Here, we ask whether deep ensembles obey a similar principle. When model complexity is properly controlled, will a single large network outperform any ensemble of smaller networks?
Deep networks in the “lazy training” regime reduce to kernel machines, with random features given by the gradients of the loss at initialization (Chizat et al., 2019). Consequently, random feature models are reliable toy models for lazy training, with the number of random features as a proxy for the number of parameters in the network (Jacot et al., 2018; Lee et al., 2019; Chizat et al., 2019; Bordelon et al., 2020; Canatar et al., 2021). For example, RFRR exhibits overfitting at finite width and sample size Atanasov et al. (2022). We therefore expect deep ensembles to obey the “no free lunch from ensembles” principle at optimal weight decay when trained in the lazy limit.
Furthermore, Theorem 2 allows us to pinpoint any violations of “no free lunch from ensembles” at optimal weight decay as a consequence of feature-learning. Feature learning can, for example, complicate the relationship between network size and performance, if the strength of feature-learning depends on network width, as in NTK parameterization (Jacot et al., 2018; Aitchison, 2020). To make fair comparisons between large models and ensembles of smaller models, we seek instead a parameterization which keeps training dynamics consistent across widths, with monotonic improvements in performance as width increases.
Maximal update parameterization ($\mu P$) (Geiger et al., 2020; Mei et al., 2018; Rotskoff & Vanden‐Eijnden, 2022; Yang & Hu, 2021; Bordelon & Pehlevan, 2022) accomplishes this desired width-consistency (Vyas et al., 2023). $\mu P$ is the unique parameterization in which the network’s infinite width limit converges and permits feature learning in finite time (Yang et al., 2022; Bordelon & Pehlevan, 2022). $\mu P$ is similar to the NTK parameterization, except we center and scale the output of the neural network inversely with a richness parameter (Chizat et al., 2019):
$\tilde{f}(x;\theta)=\frac{1}{\gamma}\Bigg{(}f(x;\theta)-f(x;\theta_{0})\Bigg{)},\qquad\gamma=\gamma_{0}\sqrt{N},\quad\eta=\eta_{0}\gamma^{2}$ (23)
so that the richness $\gamma_{0}$ and learning rate $\eta_{0}$ are constants and $\gamma$ and $\eta$ scale with network size (Geiger et al., 2020; Mei et al., 2018; Rotskoff & Vanden‐Eijnden, 2022; Yang & Hu, 2021; Bordelon & Pehlevan, 2022). At small $\gamma_{0}$, small changes in the weights are sufficient to interpolate the training data, yielding a model well-approximated as a linear model with the kernel given by the NTK at initialization. This is known as lazy learning. At large $\gamma_{0}$, large weight updates are necessary to change the network’s output, and the model learns task-relevant features. This is known as rich learning.
We train ensembles of deep convolutional neural networks (CNNs) on the CIFAR-10 image classification task, sweeping over ensemble size $K$, richness $\gamma_{0}$, and weight decay $\lambda$. We also use standard data augmentation techniques and early stopping to guard (at least partially) against over-fitting. We test a $\sim 350,000$ parameter CNN architecture with two CNN layers (figs. 5A, S8), as well as a $\sim 30$ million parameter ResNet18 architecture (figs. 5A, S9). The width of the convolutional and MLP layers are varied with $K$ to keep the total parameter count fixed (details in Appendix E). We find that in both the lazy and rich regimes, accuracy decreases monotonically with $K$, provided weight decay is tuned to its optimal value. A single exception occurs for the smaller CNN ensembles at $\gamma_{0}=10^{-0.4}$, where $K=3$ outperforms $K=2$ even at optimal weight decay. Importantly, though, monotonicity is restored when weight decay and richness $\gamma_{0}$ are jointly tuned to their optimal values (figs. 5A,B, S8, S9). Similar behavior holds when CNNs are trained without data augmentation or early stopping: provided the weight decay and richness are tuned to their optimal values, the “no free lunch from ensembles” behavior persists (fig. S10).
We also test the performance of ensembles of transformers trained on the C4 language modeling task in the online setting. After training for $5000$ steps, we find that error is monotonically increasing with $K$ across richness parameters $\gamma_{0}$ (fig. 5B).
While more comprehensive testing is required to confirm that these principles hold in general, our experiments suggest “no free lunch from ensembles” as a rule-of-thumb for deep ensembles with feature learning. In particular, we find that a single large neural network almost always outperforms an ensemble of comparable total size when $\mu P$ parameterization is used to keep training dynamics consistent across widths and either optimal weight decay or online training are used to prevent overfitting.
Figure 5: (A, B) No Free Lunch from deep CNN ensembles on CIFAR-10. At optimal weight decay, performance decays (decrease in test accuracy) monotonically with ensemble size $K$ when the total number of parameters $M$ is fixed, for lazy training ($\gamma_{0}\ll 1$) and at optimal richness. We test a small CNN architecture with 2 convolutional layers and one MLP layer (A) and ResNet18 ensembles (B) using data augmentation and optimal early stopping. (C) No Free Lunch from Transformer ensembles trained online for 5000 steps on the C4 dataset. For all $\gamma_{0}$, the performance decays (indicated by an increase in loss) monotonically with the ensemble size. Errors are optimal
@@ 7 Discussion
Another limitation of our work is the absence of feature-learning in the RFRR toy model, which prevents a direct application of our theory to deep ensembles in the rich regime. Nevertheless, we find that when deep networks are trained using $\mu$P parameterization, the “no free lunch from ensembles” principle holds empirically provided the weight decay and richness parameter $\gamma$ are tuned to their optimal values. This fact might be proven rigorously by extending a recent analytical model of feature-learning networks to the ensembled case (Bordelon et al., 2024b).
Our study also connects to recent work on scaling laws in deep learning (Kaplan et al., 2020; Hoffmann et al., 2022; Bordelon et al., 2024a; b), which observe that the test error of neural networks tends to improve predictably as a power-law with the number of parameters and the size of the dataset used during training. With our scaling-law analysis, we extend the power-laws predicted using random-feature models (Bahri et al., 2024; Maloney et al., 2022; Bordelon et al., 2020; 2024a; Defilippis et al., 2024) to the case where model size is scaled up by jointly increasing the ensemble size $K$ and parameters per ensemble member $N$ according to a “growth exponent” $\ell$ defined in eq. 21. While optimal scaling is always achieved by fixing $K$ and scaling up network size, for sufficiently easy tasks ($r>1/2$), near-optimal scaling laws can be achieved by growing both network size and ensemble size, provided network size grows quickly enough with total parameter count. Because feature-learning networks can dynamically align their representations to the target function Bordelon et al. (2024b), the scaling laws for deep ensembles may be dramatically improved by feature-learning effects.
@@ 8 Conclusion
In this work, we analyzed a trade-off between ensemble size and features-per-ensemble-member in the tractable setting of RFRR. We prove a “no free lunch from ensembles” theorem which states that optimal performance is always achieved by allocating all features to a single, large RFRR model, provided that the ridge parameter is tuned to its optimal value. A scaling-laws analysis reveals that the sharpness of this trade-off depends sensitively on the structure of the task. In particular, near-optimal scaling laws can be achieved by RFRR ensembles, provided the task is sufficiently aligned with the top modes of the limiting kernel eigenspectrum. We confirm that in deep CNN ensembles trained on a computer vision task, increasing ensemble size $K$ leads to worse performance in both a computer vision and language modeling task, provided total parameter count is fixed across $K$ values and both the weight decay and the richness parameters are tuned to their optimal values. Our results tell a story consistent with the general trend from massively ensembled predictors to large models trained by joint optimization of all parameters with respect to a single loss function. Our results have implications for model design and resource allocation in practical settings, where model size is limited.
Acknowledgments
CP is supported by NSF Award DMS-2134157, NSF CAREER Award IIS-2239780, and a Sloan Research Fellowship. WLT is supported by a Kempner Graduate Fellowship. HC was supported by the GFSD Fellowship, Harvard GSAS Prize Fellowship, and Harvard James Mills Peirce Fellowship. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence. BSR thanks Blake Bordelon, Alex Atanasov, Sabarish Sainathan, James B. Simon, and especially Jacob Zavatone-Veth for thoughtful discussion related to and comments on this manuscript.
@@ Appendix A Extending the Random-Feature Eigenframework to Ensembles
We consider an ensemble of $K$ random feature models, each with $N$ features. The total number of features is thus $M=K\cdot N$. For each model $k=1,\ldots,K$, we sample $N$ feature vectors $\{\bm{w}_{n}^{k}\}_{n=1}^{N}$ i.i.d. from a measure $\mu_{\bm{v}}$ over $\mathbb{R}^{B}$ (we will use upper indices to represent the index of the ensemble member, and lower indices to represent indices of the training examples and features). An ensemble of $K$ featurization transformations are defined as $\bm{\psi}^{k}:\bm{x}\mapsto(g(\bm{v}_{n}^{k},\bm{x}))_{n=1}^{N}$ where $g:\mathbb{R}^{B}\times\mathbb{R}^{D}\to\mathbb{R}$ is square-integrable with respect to $\mu_{\bm{x}}$ and $\mu_{\bm{v}}$. The predictions of the ridge regression models are then given as $f^{k}(\bm{x})=\bm{w}^{k}\cdot\bm{\psi}(\bm{x})$, where the weight vectors $\bm{w}^{k}$ are determined by standard linear ridge regression with a ridge parameter $\lambda$:
$\hat{\bm{w}}^{k}=\left(\frac{1}{N}{\bm{\Psi}^{k}}^{\top}\bm{\Psi}^{k}+\lambda\mathbf{I}\right)^{-1}\frac{{\bm{\Psi}^{k}}^{\top}\bm{y}}{N}$ (A.1)
Where the matrices $\bm{\Psi}^{k}\in\mathbb{R}^{N\times P}$ have columns $[\bm{\psi}^{k}(\bm{x}_{1}),\cdots,\bm{\psi}^{k}(\bm{x}_{P})]$ and the vector $\bm{y}\in\mathbb{R}^{P}$ has $[\bm{y}]_{p}=y_{p}$. For each ensemble member, this is equivalent to the kernel ridge regression predictor:
@@
Where the matrix $[\hat{\bm{H}}_{\mathcal{X}\mathcal{X}}]_{pp^{\prime}}=\hat{\bm{H}}^{k}(\bm{x}_{p},\bm{x}_{p^{\prime}})$ and the vector $[\hat{\bm{h}}_{x,\mathcal{X}}]_{p}=\hat{\bm{H}}^{k}(\bm{x},\bm{x_{p}})$ with the stochastic finite-feature kernel
@@
$E_{g}\equiv\mathbb{E}_{\bm{x}\sim\mu_{\bm{x}}}\left[(f_{\text{ens}}(\bm{x})-f_{*}(\bm{x}))^{2}\right]+\sigma_{\epsilon}^{2}$ (A.5)
@@ A.1 Spectral Decomposition of the Kernel
[…]
The singular value decomposition of $T$ ((Kato, 1966)) yields orthonormal bases $\{\zeta_{n}\}_{n=1}^{\infty}$ of $\operatorname{Ker}^{\perp}(T)\subset L^{2}(\mu_{\bm{v}})$ and $\{\phi_{n}\}_{n=1}^{\infty}$ of $L^{2}(\mu_{n})$, where $\{\eta_{t}\}_{t=1}^{\infty}$ are the eigenvalues (in decreasing order) and $\{\zeta_{t}\}_{t=1}^{\infty}$ the corresponding eigenvectors integral operator $\Sigma:L^{2}\left(\mu_{\bm{x}}\right)\rightarrow L^{2}\left(\mu_{\bm{x}}\right)$ given by
[…]

### v2

@@ 6.2 Scaling Laws for Underparameterized RFRR Ensembles
We now study the behavior of RFRR ensembles in the underparameterized regime where $N=M/K\ll P$. Here, model size $N$ is the “bottleneck” that dictates the scaling behavior of the risk (Bahri et al., 2024; Maloney et al., 2022; Bordelon et al., 2024a; Atanasov et al., 2024). In particular, we ask how $E_{g}^{K}$ scales with total model size $M$ when $K$ and $N$ scale jointly with $M$ according to a “growth exponent” $\ell$:
@@
so that when $\ell=0$ the ensemble grows with $M$ by adding additional ensemble members of a fixed size, and when $\ell=1$ the ensemble grows by adding parameters to a fixed number of networks (see fig. 3). Under the standard “source” and “capacity” constraints on the task eigenstructure (Cui et al., 2023; Caponnetto & De Vito, 2006; Bordelon et al., 2020; Defilippis et al., 2024) (the kernel eigenspectrum decays as $\eta_{t}\sim t^{-\alpha}$ with $\alpha>1$ and the target weights decay as $\bar{w}_{t}^{2}\eta_{t}\sim t^{-(1+2\alpha r)}$), we find:
$E_{g}^{K}\sim M^{-s}$ (26)
$s=\min\left(2\alpha\ell\min(r,1),1-\ell+2\alpha\ell\min\left(r,\frac{1}{2}\right)\right)\,$, (27)
with $s=2\alpha\ell\min(r,1)$ corresponding to the scaling of the bias and $s=1-\ell+2\alpha\ell\min\left(r,1/2\right)$ corresponding the scaling of the variance (reduced by a factor of $1/K$). A full derivation is provided in Appendix D. These results reify the “no free lunch” result, as the optimal scaling law is always achieved when $\ell=1$. For difficult tasks, defined as having $r<1/2$, bias always dominates the error scaling and the scaling exponent increases linearly with $\ell$. However, when $r>1/2$, there will be a certain value $\ell^{*}$ above which error scaling is dominated by the variance term. When $r>1/2$, the scaling exponent of the variance increases from $1$ to $\alpha$ over the range $\ell\in[0,1]$. If $\alpha\gtrsim 1$, this can approach a flat line, and the dependence of the scaling exponent on $\ell$ can become weak, so that near-optimal scaling can be achieved for any $\ell>\ell^{*}$. When $1/2<r<1$, this transition occurs at $\ell^{*}=1/(1+\alpha(2r-1))$ and when $r>1$ it occurs at $\ell^{*}=1/(1+\alpha)$.
We plot these scaling laws with $\alpha=1.5$ in the regimes where $r<1/2$, $1/2<r<1$, and $r>1$ in fig. 4, along with the results of numerical simulations of linear RF regression on synthetic Gaussian datasets. As anticipated, for difficult tasks, where $r<1/2$, the scaling law improves linearly with $\ell$. However, for easier tasks ($r>1/2$), we see that, a near-optimal scaling law can be achieved as long as $\ell>\ell^{*}$ (fig. 4, center and right columns).
@@
Figure 5: Scaling laws provide a good description of width-bottlenecked RFRR ensembles.(A) we plot error as a function of $M$ at optimal ridge value for $\operatorname{ReLU}$ random-feature models applied to the binarized CIFAR-10 (left) and MNIST (right) classification tasks. (B) We plot theoretically predicted scaling exponents (eq. 26) for the bias and variance contributions to risk, as well as empirical power-law fits to risk in numerical simulations of RFRR models (see Appendix E.3, fig. S7)
@@ 7 Discussion
Our results show that in RFRR ensembles any violations of the “more is better” or “no free lunch from ensembles” principles are a result of overfitting at non-optimal ridge. This suggests that violations of these principles in other model types may also be a result of overfitting. Because deep networks reduce to kernel methods in the lazy training limit (Chizat et al., 2019; Jacot et al., 2018), we may pinpoint any violations of these principles in deep ensembles as a consequence specifically of feature learning. Future work should extend recent analytical treatments of feature-learning networks (Bordelon et al., 2024b) to the ensembled case.
Our results relate to a line of work on “sketched” or subsampled ridge ensembles, which prove that an infinite ensemble of regression models trained on random projections of the data achieve equivalent test risk to a single regression model trained on the full data, with an increased ridge parameter (Patil & LeJeune, 2024; Atanasov et al., 2024; LeJeune et al., 2020; Patil et al., 2023). Translated to the kernel eigenframework under Gaussian universality, these results imply that in the limit $K\to\infty$, RFRR ensembles converge on the same test risk as the limiting (infinite-feature) kernel predictor. It follows that no ensemble of RFRR models can outperform the limiting kernel predictor at optimal ridge, corresponding to the $M\to\infty$ limit of theorem 5.1 with $K^{\prime}=1$. Because it applies also at finite $M$ and $K^{\prime}>1$, theorem 5.1 is a stronger statement.
Our study also connects to recent work on scaling laws in deep learning (Kaplan et al., 2020; Hoffmann et al., 2022; Bordelon et al., 2024a, b), which observe that the test error of neural networks tends to improve predictably as a power-law with the number of parameters and the size of the dataset used during training. With our scaling-law analysis, we extend the power-laws predicted using random-feature models (Bahri et al., 2024; Maloney et al., 2022; Bordelon et al., 2020, 2024a; Defilippis et al., 2024) to the case where model size $N$ and ensemble size $K$ are scaled jointly according to a novel “growth exponent” $\ell$ defined in eq. 25. While optimal scaling is always achieved by fixing $K$ and scaling $N$, for sufficiently easy tasks ($r>1/2$), near-optimal scaling laws can be achieved by growing both $N$ and $K$, provided $N$ grows quickly enough with total parameter count. Because feature-learning networks can dynamically align their representations to the target function (Bordelon et al., 2024b), the scaling laws for deep ensembles may be dramatically improved by feature-learning effects.
@@ 8 Conclusion
In this work, we analyzed a trade-off between ensemble size and features-per-ensemble-member in the tractable setting of RFRR. We prove a “no free lunch from ensembles” theorem which states that optimal performance is always achieved by allocating all features to a single, large RFRR model, provided that the ridge parameter is tuned to its optimal value. A scaling-laws analysis reveals that the sharpness of this trade-off depends sensitively on the structure of the task. In particular, near-optimal scaling laws can be achieved by RFRR ensembles, provided the task is sufficiently aligned with the top modes of the limiting kernel eigenspectrum. Our results tell a story consistent with the general trend from massively ensembled predictors to large models trained by joint optimization of all parameters with respect to a single loss function. Our results have implications for model design and resource allocation in practical settings, where model size is limited.
Acknowledgements
C.P. is supported by NSF grant DMS-2134157, NSF CAREER Award IIS-2239780, DARPA grant DIAL-FP-038, a Sloan Research Fellowship, and The William F. Milton Fund from Harvard University. WLT is supported by a Kempner Graduate Fellowship. HC was supported by the GFSD Fellowship, Harvard GSAS Prize Fellowship, and Harvard James Mills Peirce Fellowship. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence. BSR thanks Blake Bordelon, Alex Atanasov, Sabarish Sainathan, James B. Simon, and especially Jacob Zavatone-Veth for thoughtful discussion related to and comments on this manuscript.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Extending the Random-Feature Eigenframework to Ensembles
We consider an ensemble of $K$ random feature models, each with $N$ features. The total number of features is thus $M=K\cdot N$. For each model $k=1,\ldots,K$, we sample $N$ feature vectors $\{\bm{w}_{n}^{k}\}_{n=1}^{N}$ i.i.d. from a measure $\mu_{\bm{v}}$ over $\mathbb{R}^{B}$ (we will use upper indices to represent the index of the ensemble member, and lower indices to represent indices of the training examples and features). An ensemble of $K$ featurization transformations are defined as $\bm{\psi}^{k}:\bm{x}\mapsto(g(\bm{v}_{n}^{k},\bm{x}))_{n=1}^{N}$ where $g:\mathbb{R}^{B}\times\mathbb{R}^{D}\to\mathbb{R}$ is square-integrable with respect to $\mu_{\bm{x}}$ and $\mu_{\bm{v}}$. The predictions of the ridge regression models are then given as $f^{k}(\bm{x})=\bm{w}^{k}\cdot\bm{\psi}(\bm{x})/\sqrt{N}$, where the weight vectors $\bm{w}^{k}$ are determined by standard linear ridge regression with a ridge parameter $\lambda$:
$\hat{\bm{w}}^{k}=\left(\frac{1}{N}{\bm{\Psi}^{k}}^{\top}\bm{\Psi}^{k}+\lambda\mathbf{I}\right)^{-1}\frac{{\bm{\Psi}^{k}}^{\top}\bm{y}}{\sqrt{N}}$ (A.1)
Where the matrices $\bm{\Psi}^{k}\in\mathbb{R}^{N\times P}$ have columns $\bm{\psi}^{k}(\bm{x}_{1}),\cdots,\bm{\psi}^{k}(\bm{x}_{P})]$ and the vector $\bm{y}\in\mathbb{R}^{P}$ has $\bm{y}]_{p}=y_{p}$. For each ensemble member, this is equivalent to the kernel ridge regression predictor:
@@
Where the matrix $\hat{\bm{H}}_{\mathcal{X}\mathcal{X}}]_{pp^{\prime}}=\hat{\bm{H}}^{k}(\bm{x}_{p},\bm{x}_{p^{\prime}})$ and the vector $\hat{\bm{h}}_{x,\mathcal{X}}]_{p}=\hat{\bm{H}}^{k}(\bm{x},\bm{x_{p}})$ with the stochastic finite-feature kernel
@@
$E_{g}^{K}\equiv\mathbb{E}_{\bm{x}\sim\mu_{\bm{x}}}\left[(f_{\text{ens}}(\bm{x})-f_{*}(\bm{x}))^{2}\right]+\sigma_{\epsilon}^{2}$ (A.5)
@@ A.1 Spectral Decomposition of the Kernel
$(Tr)(x)=\int_{\mathbb{R}^{B}}r(\bm{v})g(\bm{v},\bm{x})d\mu_{\bm{v}}(\bm{v})$ (A.8)
The singular value decomposition of $T$ (Kato, 1966) yields orthonormal bases $\{\zeta_{n}\}_{n=1}^{\infty}$ of $\operatorname{Ker}^{\perp}(T)\subset L^{2}(\mu_{\bm{v}})$ and $\{\phi_{n}\}_{n=1}^{\infty}$ of $L^{2}(\mu_{n})$, where $\{\eta_{t}\}_{t=1}^{\infty}$ are the eigenvalues (in decreasing order) and $\{\zeta_{t}\}_{t=1}^{\infty}$ the corresponding eigenvectors integral operator $\Sigma:L^{2}\left(\mu_{\bm{x}}\right)\rightarrow L^{2}\left(\mu_{\bm{x}}\right)$ given by
$(\Sigma u)(\bm{x})=\int_{\mathbb{R}^{D}}u(\bm{x}^{\prime})\bm{H}(\bm{x}^{\prime},\bm{x})d\mu_{\bm{x}}(\bm{x}^{\prime})$ (A.9)